=== 2019/3/11(Mon.) ===

=== Schedule ===
 * Center
  * --(2/26-27: Venting)--
  * 2/28-3/15: Health check and maintenance for VIS
   * --(PRM: One flag was dropped. -> leave it)--
   * --(3/6 PR3: range -> incline by 20-30mrad at mirror because of bistable.)--
   * 3/4-8 PRM: Misalignment does not reach where beam hits damper.
   * ~3/15 BS:  Yaw offset adjustment by hand. -> 3/5-6

  * --(3/5 afternoon or 3/6 morning: Coolers for ITMs will be stopped for 2-3 hours to change water.)--
  * 3/18-19: Re-alignment
  * 3/19: Closing
  * 3/20: pump down

 * Yend
  * --(1/25-2/20: Yend Pcal finalization)-- 
  * --(3/5: EYA remaining works by Enomoto, Inoue.)--
  * --(2/25-3/5: Yend WAB finalization)--
  * --(3/6: Y end closing (no laser hazard at least in the morning))--
  * --(3/7: Y end pump down)--
   * --(3/7,8: leak test)--

 * Xend
  * --(2/11-2/14: Xend NAB installation)--
  * --(2/21: Removing FC, Payload release)--
   * --(2/27: ETMX release)--
  * --(2/23-3/6: Xend Pcal finalization 1)--
   * --(Periscope installation)--
   * --(3/6: Beam position by IR and Green)--
  * --(3/5 AOS: EY and EX(if they have time) work.)--
  * --(3/6: Green X alignment (no IR), and no hazard at Y end.)--
  * 3/11, 18,19: Xend WAB finalization.
  * 3/11~3/15: ETMX tuning.
  * 3/19: Heatlink etc.
  * 3/19: ETMX finalization by Takahashi
  * 3/18,19: ETMX alignement
  * 3/21 closing
  * 3/22 Pumping down
   * Leak test
 
 * 3/18~ : Night work starts, daily meeting will be moved to 1PM
  * Summary of work will last for a while in the morning meeting.

 * 2/18-2/22: PRMI commissioning
 * 3/25~?: DRMI commissioning
 * 3/25-4/19: Y arm commissioning
 * 3/25-4/19: FPMI commissioning

==== Note ====
 * Night work does not start from 3/11 due to the 2weeks delay by VIS troubles.
  * It does not mean that night work is prohibited but it means that officially night shift responsible person won't be assigned and we have no daily meeting at 1pm.
  * Official night work starts 3/18 or 3/25, but not determined yet.

=== Weekend ===

==== Report of work ====
 * Type Bp: health check PR3 seems to be OK.
 * Type B: SR2 and SR3 SRM are fine. SRM had glitch.
  * IM drift seem to be zero and opposite direction.

 * Type A: F3 of ETMY seems to be stuck somehow, by that temperature changed or vacuum started?
  * Temperature is low under 20degree at 1st floor of ETMY.

 * PEM: finished work at ETMX area.


==== Work plan ====
 * AOS: Finalization of WAB at ETMX will start today. It lasts this week.